DIRECTIONS: Solve for the z-computed value of the following. Write your answer to the

nearest hundredths. Show the complete solution.

1. x̅ = 9.2 μ = 10 σ =3 n = 68

2. x̅ = 28.3 μ = 26 σ = 4.5 n = 80

3. x̅ = 72.2 μ = 75 σ = 5.8 n = 118

4. x̅ = 49.6 μ = 52 σ = 7 n = 160

5. x̅ = 92 μ = 100 σ = 12 n = 130


1
Expert's answer
2022-04-26T13:23:31-0400

"Z=\frac{x-\mu}{\sigma/\sqrt{n}}"

1."Z=\frac{9.2-10}{3/\sqrt{68}}=-2.20"

2."Z=\frac{28.3-26}{4.5/\sqrt{80}}=4.57"

3. "Z=\frac{72.2-75}{5.8/\sqrt{118}}=-5.24"

4. "Z=\frac{49.6-52}{7/\sqrt{160}}=-4.34"

5."Z=\frac{92-100}{12/\sqrt{130}}=-7.6"